OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=8, (2,1) the equation of the line is y=?

OpenStudy (sidsiddhartha):

equation of a line having slope=m and passing through points (x1,y1) is \[(y-y_1)=m(x-x_1)\] here you have m=8 x1=2 and y1=1 now just plug the values

OpenStudy (anonymous):

How do i do that?

OpenStudy (sidsiddhartha):

m=8 x1=2 and y1=1 just put them in the above equation

OpenStudy (anonymous):

1y-1y=8x-x

OpenStudy (sidsiddhartha):

nope its \[y-y_1=m(x-x_1)\\y-1=8(x-2)\] ok?

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

:)

OpenStudy (anonymous):

thats the answer?

OpenStudy (sidsiddhartha):

u can simplify it a bit use simple alzebra

OpenStudy (anonymous):

\[y=8 x-15 \]

OpenStudy (sidsiddhartha):

yup

OpenStudy (anonymous):

its saying thats wrong @robtobey

OpenStudy (mrnood):

@MarthaHearst @robtobey - it is is not encouraged to post the complete answer. The OP had plenty of opportunity to do some of this with help above. The answer IS correct - your book must be wrong, or you have posted the question incorrectly

OpenStudy (anonymous):

@MrNood who are you?

OpenStudy (anonymous):

apply the formula y2-y1=m(x2-x1) where m is the slope

OpenStudy (mrnood):

@rosebird Why would she need to do that now? She has been given the correct answer by someone else, and has had no need to begin to understand the straight line formula concepts that she is studying

OpenStudy (anonymous):

ok mind ur own busines please @MrNood

OpenStudy (mrnood):

It is everyone's business that education gets pupils to understand, NOT just to simply tick a box on a multichoice question

OpenStudy (anonymous):

yeh i know but i posted the answer so whats the matter with u?

OpenStudy (anonymous):

@MrNood